OpenStudy (anonymous):

what is the volume of a spere with a diameter of 24in.

OpenStudy (hunus):

\[\Large V= \frac{4}{3} \pi r^3\] The radius is half of the diameter so r will be 24/2= 12

OpenStudy (anonymous):

thx dude is that in term of pi

OpenStudy (hunus):

Yes, just don't it will be (4/3)r^3 * pi

OpenStudy (anonymous):

ok thx